# Autonomous Stabilization of Floquet States Using Static Dissipation

### v1

@@ Acknowledgements.
We thank Maya Amouzegar for assistance with device fabrication, Benjamin Cochran for contributions to device modeling, IBK Adisa for assistance with early measurements; and Chris Laumann and Vedika Khemani for helpful comments. During completion of this manuscript, we became aware of an unrelated proposal [71] in which the ability to achieve resonance in a Jaynes-Cummings model using strong Rabi drives was harnessed for a topological pump. This work was supported by ARL (Grants no. W911NF-19-2-0181 and W911NF-17-S-0003), the University of Maryland, and AFOSR (Grants No. FA9550-20-1-0235 and FA9550-24-1-0121). MR received support from the National Science Foundation (PFC at JQI Grant No. PHY-1430094), the Laboratory for Physicsal Sciences graduate fellowship, and ARCS. QY received support from AFOSR Grant No. FA9550-21-1-0129. DL was supported by the Laboratory for Physical Sciences, a Stanford Q-FARM Bloch Postdoctoral Fellowship, and the Packard Foundation through a Packard Fellowship in Science and Engineering.
@@ Appendix B Synthesizing a Rotating Field
using a microwave drive ($X$-drive) to provide the $x$ component and frequency modulation of the qubit ($Z$-drive) to provide the $z$ component of the field. Here we sketch the general theoretical framework for producing an effective magnetic field of this form for a transmon qubit, which does not natively exhibit a magnetic moment. Technical details of the required control signals and calibration protocols are described in [72].
The transition frequency of a transmon formed using a DC SQUID in parallel with a capacitor is controlled by the external flux threaded through the SQUID [57, 58]. The $z$-component of the desired drive is therefore relatively straightforward to implement. A DC bias, set by an external magnet to reduce noise, is used to set the desired average qubit frequency, and an AC bias is applied using an on-chip flux-bias line to modulate the qubit frequency about this value:
@@
where $\omega_{q_{0}}$ is the qubit frequency set by the DC bias, and $B_{0}$, $\omega_{\mathrm{mod}}$ are the amplitude and rotation frequency of the applied magnetic field respectively. The DC bias point is maintained by an active servo [72] to keep the average qubit frequency at the desired point even in the presence of environmental drift. The AC bias line is configured with a $5$ MHz low-pass filter at the base-plate of the dilution refrigerator to cut out high-frequency noise. The required AC current amplitude is calibrated by performing qubit spectroscopy after applying a flux bias pulse. The group delay induced by the filter and line is compensated by adding a calibrated buffer time that matches the time between sending flux bias pulse and observing a response in the qubit spectroscopy. See Ref. [72] for details.
In contrast, the $x$-component of the desired effective magnetic field is not realized by applying a laboratory-frame magnetic field. Instead, we make use of the well-known fact that driving a two-level system with an oscillatory electric field leads to Rabi oscillations between the |$\ket{g}$ and |$\ket{e}$ states, which can be mapped to the dynamics of spin in a magnetic field [59]. The bare laboratory-frame Hamiltonian takes the from
@@
where $\Delta_{d}$ = $\omega_{q}-\omega_{d}$ is the detuning between the qubit and the drive, and $\Omega$ is the Rabi rate [59].
@@
A second important hardware constraint to note is that the sign of the microwave drive must be reversed during the cycle in order to produce negative values of $B_{x}$. As a result, a conventional amplitude modulator is not sufficient to realize the required field. We make use of a vector microwave generator with an IQ modulator in order to have control over both the amplitude and the sign of the microwave drive [72, 58]. We calibrate our $X$-field using standard Rabi oscillation measurements by varying the amplitude of the applied microwave drive and fitting the resulting oscillation frequencies.
@@ Appendix D Numerical simulations of the augmented spin-cavity and boosting setups
with $L_{1}=\sqrt{\kappa}a$ (the jump operator modelling photon loss), $L_{2}=\sqrt{\Gamma}\sigma^{-}$ (the jump operator modelling qubit loss, or $T_{1}$ relaxation), and $\{A,B\}=AB+BA$ being the anticommutator. In Sec. 5 we use a similar Lindblad framework in a Hilbert space with two cavities, but with $H_{B}^{+}$ replaced by $H_{\mathrm{boost}}$ and a single jump operator $L_{1}=\sqrt{\kappa_{\mathrm{s}}}a_{\mathrm{s}}$. We solve for $\rho(t)$ given some initial conditions by direct integration in a truncated cavity Hilbert space (we have checked that the state remains far from the truncation throughout the evolution), using the master equation methods of QuTiP [73]. The data displayed in Figs. 5, 9, and 11 is obtained in this way.
@@
In the Sec V, it is claimed that the stabilized protocol resulted in a boosted state with lower variance. This is visible in Fig. 9, which shows the same data on a larger scale. At, for instance, $t^{*}=12T_{\mathrm{mod}}$, the unstabilized protocol produces a state with modal value $n_{\mathrm{b}}=22$, as does the stabilized protocol. However, the unstabilized state has a standard deviation in $n_{\mathrm{b}}$ at this time of $\mathrm{Std}[n_{\mathrm{b}}]\approx 5.98$—more than twice the value of the stabilized state, which has $\mathrm{Std}[n_{\mathrm{b}}]\approx 2.72$. The latter is less than the standard deviation of a coherent state with a mean of 22 photons, $\sqrt{22}\approx 4.69$, so that the stabilized state can outperform a coherent state of the same size in phase estimation measurements [60, 61].
@@ Appendix F Stabilization Rate and Fidelity
The theoretical curves come from a simulation of the Lindblad master equation [73] described in Appendix D, with jump operators for qubit depolarization, qubit dephasing, and cavity decay. The Hamiltonian parameters of the numerical simulation are set to the experimentally calibrated energy-scales shown in Table 1. The strengths of the cavity and qubit decay jump operators are set to $\kappa=1/T_{\mathrm{cav}}$ and $\Gamma_{q}=1/T_{1}$, using the experimentally measured values shown in Appendix G.
@@ Appendix G Qubit characterization
We perform standard characterization of the various decoherence timescales in the qubit-cavity system with the qubit at $4.9$ GHz, far detuned from the main cavity. Corresponding measurements are shown in Fig. 12. We extract the qubit decay time $T_{1}=11.5\ \mu$s from an exponential fit to the data after initializing the qubit in the |$\ket{e}$ and waiting a time $\tau$, with the qubit far detuned from both main cavity. Purcell loss from the main cavity provides a negligible contribution to measurement, so $T_{1}=11.5\ \mu$s, therefore, represents the decay rate of the qubit into channels other than the main cavity.
To find pure dephasing time $T_{\phi}=570$ ns, we perform a Ramsey-style measurement where two $\pi/2$ pulses are applied to the qubit with a variable wait time. However, to maintain a large separation of scales between the oscillatory and decaying time constants of the Ramsey signal, we rotate the phase of the second pulse at $5$ MHz to apply a virtual detuning between the drive and qubit without loss of contrast. We use a gaussian envelope to compute the dephasing time [58] instead of the regular exponential envelope, due to the low-frequency noise present in tunable frequency qubits away from a sweet spot. We extract $T_{\mathrm{echo}}=3\ \mu$s from the standard Hahn echo sequence and fit an exponential model to the data.
@@
Figure 12: System coherence times. (a) |$\ket{e}$ state population as a function of delay time after qubit excitation. An exponential fit gives $T_{1}=11.5\ \mu$s. (b) Ramsey oscillations using a $5$ MHz virtual detuning on the second $\pi/2$-pulse to extract $T_{\phi}=570$ ns. (c) |$\ket{e}$-state population after a standard Hahn echo sequence, giving $T_{\mathrm{echo}}=3\ \mu$s. The six-fold increase between $T_{\phi}$ and $T_{\mathrm{echo}}$ indicates the presence of a large low-frequency noise component in the qubit frequency. (d) Heterodyne measurement of the main cavity field after populating it with a coherent state. The field decay time $T_{\mathrm{cav}}=3.8\ \mu$s is extracted from a fit to a damped oscillatory response (envelope of the fit is indicated by dashed lines). The photon lifetime corresponds to $T_{\mathrm{cav}}/2$. Presented data is the average of $10$ repeated traces, each consisting of $10^{4}$ shots per time point. Error bars are computed as the standard error of the mean of the repeated traces. The error bars are smaller than the marker size for panels (a), (b), (c) and not shown for panel (d).
@@ Appendix H Experimental error bars and qubit state readout
The qubit state is read out using standard two-tone pulsed spectroscopy on the readout cavity [57, 58]. In this method, |$\ket{g}$ and |$\ket{e}$ correspond to two different resonant frequencies of the readout cavity. The qubit state can be inferred by measuring the amplitude and phase of a microwave pulse at the cavity frequency, which can be converted to IQ points (the cartesian representation of the sine and cosine components of the cavity response). By taking reference |$\ket{g}$ and |$\ket{e}$ measurements, we can convert an output voltage to a normalized qubit population:

### v2

@@ Acknowledgements.
We thank Maya Amouzegar for assistance with device fabrication, Benjamin Cochran for contributions to device modeling, IBK Adisa for assistance with early measurements; and Vedika Khemani, Chris Laumann, and Paul Schindler for helpful comments. During completion of this manuscript, we became aware of an unrelated proposal Lin et al. (2024) in which the ability to achieve resonance in a Jaynes-Cummings model using strong Rabi drives was harnessed for a topological pump. This work was supported by ARL (Grants no. W911NF-19-2-0181 and W911NF-17-S-0003), the University of Maryland, NSF (QLCI grant OMA-2120757), and AFOSR (Grants No. FA9550-20-1-0235 and FA9550-24-1-0121). MR received support from the National Science Foundation (PFC at JQI Grant No. PHY-1430094), the Laboratory for Physical Sciences graduate fellowship, and ARCS. QY received support from AFOSR Grant No. FA9550-21-1-0129. DL was supported by the Laboratory for Physical Sciences, a Stanford Q-FARM Bloch Postdoctoral Fellowship, and the Packard Foundation through a Packard Fellowship in Science and Engineering.
@@ Appendix B Synthesizing a Rotating Field
using a microwave drive ($X$-drive) to provide the $x$ component and frequency modulation of the qubit ($Z$-drive) to provide the $z$ component of the field. Here we sketch the general theoretical framework for producing an effective magnetic field of this form for a transmon qubit, which does not natively exhibit a magnetic moment. Technical details of the required control signals and calibration protocols are described in Ritter et al. (2025).
The transition frequency of a transmon formed using a DC SQUID in parallel with a capacitor is controlled by the external flux threaded through the SQUID Blais et al. (2021); Krantz et al. (2019). The $z$-component of the desired drive is therefore relatively straightforward to implement. A DC bias, set by an external magnet to reduce noise, is used to set the desired average qubit frequency, and an AC bias is applied using an on-chip flux-bias line to modulate the qubit frequency about this value:
@@
where $\omega_{q_{0}}$ is the qubit frequency set by the DC bias, and $B_{0}$, $\omega_{\mathrm{mod}}$ are the amplitude and rotation frequency of the applied magnetic field respectively. The DC bias point is maintained by an active servo Ritter et al. (2025) to keep the average qubit frequency at the desired point even in the presence of environmental drift. The AC bias line is configured with a $5$ MHz low-pass filter at the base-plate of the dilution refrigerator to cut out high-frequency noise. The required AC current amplitude is calibrated by performing qubit spectroscopy after applying a flux bias pulse. The group delay induced by the filter and line is compensated by adding a calibrated buffer time that matches the time between sending flux bias pulse and observing a response in the qubit spectroscopy. See Ref. Ritter et al. (2025) for details.
In contrast, the $x$-component of the desired effective magnetic field is not realized by applying a laboratory-frame magnetic field. Instead, we make use of the well-known fact that driving a two-level system with an oscillatory electric field leads to Rabi oscillations between the |$\ket{g}$ and |$\ket{e}$ states, which can be mapped to the dynamics of spin in a magnetic field Walls and Milburn (2008). The bare laboratory-frame Hamiltonian takes the from
@@
where $\Delta_{d}$ = $\omega_{q}-\omega_{d}$ is the detuning between the qubit and the drive, and $\Omega$ is the Rabi rate Walls and Milburn (2008).
@@
A second important hardware constraint to note is that the sign of the microwave drive must be reversed during the cycle in order to produce negative values of $B_{x}$. As a result, a conventional amplitude modulator is not sufficient to realize the required field. We make use of a vector microwave generator with an IQ modulator in order to have control over both the amplitude and the sign of the microwave drive Ritter et al. (2025); Krantz et al. (2019). We calibrate our $X$-field using standard Rabi oscillation measurements by varying the amplitude of the applied microwave drive and fitting the resulting oscillation frequencies.
@@ Appendix D Numerical simulations of the augmented spin-cavity and boosting setups
with $L_{1}=\sqrt{\kappa}a$ (the jump operator modelling photon loss), $L_{2}=\sqrt{\Gamma}\sigma^{-}$ (the jump operator modelling qubit loss, or $T_{1}$ relaxation), and $\{A,B\}=AB+BA$ being the anticommutator. In Sec. 5 we use a similar Lindblad framework in a Hilbert space with two cavities, but with $H_{B}^{+}$ replaced by $H_{\mathrm{boost}}$ and a single jump operator $L_{1}=\sqrt{\kappa_{\mathrm{s}}}a_{\mathrm{s}}$. We solve for $\rho(t)$ given some initial conditions by direct integration in a truncated cavity Hilbert space (we have checked that the state remains far from the truncation throughout the evolution), using the master equation methods of QuTiP Johansson et al. (2013). The data displayed in Figs. 5, 9, and 11 is obtained in this way.
@@
In the Sec V, it is claimed that the stabilized protocol resulted in a boosted state with lower variance. This is visible in Fig. 9, which shows the same data on a larger scale. At, for instance, $t^{*}=12T_{\mathrm{mod}}$, the unstabilized protocol produces a state with modal value $n_{\mathrm{b}}=22$, as does the stabilized protocol. However, the unstabilized state has a standard deviation in $n_{\mathrm{b}}$ at this time of $\mathrm{Std}[n_{\mathrm{b}}]\approx 5.98$—more than twice the value of the stabilized state, which has $\mathrm{Std}[n_{\mathrm{b}}]\approx 2.72$. The latter is less than the standard deviation of a coherent state with a mean of 22 photons, $\sqrt{22}\approx 4.69$, so that the stabilized state can outperform a coherent state of the same size in phase estimation measurements Giovannetti et al. (2011); Tóth and Apellaniz (2014).
@@ Appendix F Stabilization Rate and Fidelity
The theoretical curves come from a simulation of the Lindblad master equation Johansson et al. (2013) described in Appendix D, with jump operators for qubit depolarization, qubit dephasing, and cavity decay. The Hamiltonian parameters of the numerical simulation are set to the experimentally calibrated energy-scales shown in Table 1. The strengths of the cavity and qubit decay jump operators are set to $\kappa=1/T_{\mathrm{cav}}$ and $\Gamma_{q}=1/T_{1}$, using the experimentally measured values shown in Appendix G.
@@ Appendix G Qubit characterization
We perform standard characterization of the various decoherence timescales in the qubit-cavity system with the qubit at $4.9$ GHz, far detuned from the main cavity. Corresponding measurements are shown in Fig. 12. We extract the qubit decay time $T_{1}=11.4\ \mu$s from an exponential fit to the data after initializing the qubit in the |$\ket{e}$ and waiting a time $\tau$, with the qubit far detuned from both main cavity. Purcell loss from the main cavity provides a negligible contribution to measurement, so $T_{1}=11.4\ \mu$s, therefore, represents the decay rate of the qubit into channels other than the main cavity.
To find pure dephasing time $T_{\phi}=590$ ns, we perform a Ramsey-style measurement where two $\pi/2$ pulses are applied to the qubit with a variable wait time. However, to maintain a large separation of scales between the oscillatory and decaying time constants of the Ramsey signal, we rotate the phase of the second pulse at $5$ MHz to apply a virtual detuning between the drive and qubit without loss of contrast. We use a gaussian envelope to compute the dephasing time Krantz et al. (2019) instead of the regular exponential envelope, due to the low-frequency noise present in tunable frequency qubits away from a sweet spot. We extract $T_{\mathrm{echo}}=3\ \mu$s from the standard Hahn echo sequence and fit an exponential model to the data.
@@
Figure 12: System coherence times. (a) |$\ket{e}$ state population as a function of delay time after qubit excitation. An exponential fit gives $T_{1}=11.4\ \mu$s. (b) Ramsey oscillations using a $5$ MHz virtual detuning on the second $\pi/2$-pulse to extract $T_{\phi}=590$ ns. (c) |$\ket{e}$-state population after a standard Hahn echo sequence, giving $T_{\mathrm{echo}}=3\ \mu$s. The six-fold increase between $T_{\phi}$ and $T_{\mathrm{echo}}$ indicates the presence of a large low-frequency noise component in the qubit frequency. (d) Heterodyne measurement of the main cavity field after populating it with a coherent state. The field decay time $T_{\mathrm{cav}}=3.8\ \mu$s is extracted from a fit to a damped oscillatory response (envelope of the fit is indicated by dashed lines). The photon lifetime corresponds to $T_{\mathrm{cav}}/2$. Presented data is the average of $10$ repeated traces, each consisting of $10^{4}$ shots per time point. Error bars are computed as the standard error of the mean of the repeated traces. The error bars are smaller than the marker size for panels (a), (b), (c) and not shown for panel (d).
@@ Appendix H Experimental error bars and qubit state readout
The qubit state is read out using standard two-tone pulsed spectroscopy on the readout cavity Blais et al. (2021); Krantz et al. (2019). In this method, |$\ket{g}$ and |$\ket{e}$ correspond to two different resonant frequencies of the readout cavity. The qubit state can be inferred by measuring the amplitude and phase of a microwave pulse at the cavity frequency, which can be converted to IQ points (the cartesian representation of the sine and cosine components of the cavity response). By taking reference |$\ket{g}$ and |$\ket{e}$ measurements, we can convert an output voltage to a normalized qubit population: